# Fault-tolerant hyperbolic Floquet quantum error correcting codes

### v3

@@ Appendix A Details of Error Detection and Correction
In a measurement error, the outcome of a check’s measurement is faulty. This faulty outcome is then utilized in inferring two syndromes, causing them to exhibit detection events. An example is illustrated in Fig. 9(a). In this example, the highlighted green check’s outcome at round ${r=3n}$ is erroneous, while all other measurements are accurate. This faulty green check leads to the blue plaquette syndrome at round ${r=3n}$ and the red plaquette syndrome at round ${r=3n+1}$ exhibiting detection events. Given the assumption that all other measurements are accurate, the next measurement of the same green check at round ${r=3n+3}$ relates to the accurate value of the plaquette syndromes. Thus, the same blue and red syndromes generate additional detection events three rounds later. A measurement error, therefore, is linked to four detection events, forming a hyper-edge in the syndrome lattice, depicted with a dashed parallelogram in Fig. 9(a).
@@
Figure 9: Detection events of a measurement error. (a) The faulty measurement outcome of the highlighted green check at round ${r=3n}$ is used to infer an incorrect value for the adjacent blue and red plaquette syndromes at rounds ${r=3n}$ and ${r=3n+1}$. Later correct measurement of the same check leads to correct inference of these syndromes three rounds later. In both cases, the inferred syndrome differs from the most recent inference and gives rise to a detection event. Hence, a measurement error has four detection events forming a hyper-edge in the syndrome lattice, shown with a dashed parallelogram. (b) A combination of three simple Pauli errors on one of the qubits of the faulty check, denoted with a circle, gives rise to the same detection events. These equivalent Pauli errors have the same effects on the logical operators as the measurement error.
@@
Consider the pair of detection events linked to a simple Pauli error. There are four different Pauli errors that could trigger these two detection events and, thus, are indistinguishable from one another. Without any loss of generality, assume the simple Pauli error is an X error. The two plaquette syndromes with detection events correspond to adjacent blue and red octagons [see Fig. 8(b)]. These two octagons identify the location of the physical qubit affected by the X error to their shared green edge, without distinguishing which of the two ends of that edge it is. We label these two qubits as $q_{1}$ and $q_{2}$. Moreover, the X error could have occurred either between the most recent XX and YY measurements or between the most recent ZZ and XX measurements. These possibilities for error’s support qubit, $q_{1}$ or $q_{2}$, and its timing, between ZZ and XX measurements or between XX and YY measurements, give rise to the four indistinguishable errors with identical detection events.
@@
Now consider a logical operator, $L$, that contains one of these two qubits ($q_{1}$, without loss of generality) in its closed path. Between the most recent ZZ and XX measurements, $L$’s expression can either be type one (Z Paulis on green checks) or type two (X Paulis on red checks). In each case, between the most recent XX and YY measurements, $L$’s expression will be type two (Y Paulis on green checks) or type one (X Paulis on blue checks), respectively. The X error on $q_{1}$ affects $L$ only if $L$’s expression features Y or Z Paulis on $q_{1}$ at the time of the error. This holds true if $L$ is a type one and then a type two logical operator, in addition to having the green edge of $q_{1}$’s vertex along its closed path. In this case, $L$’s expression during the two possible time intervals includes Y or Z Paulis on both $q_{1}$ and $q_{2}$. Therefore, all four indistinguishable simple Pauli errors impact logical operators in the same way, and their indistinguishability does not pose an obstruction in correcting their effects on logical operators.
@@
A compound Pauli error is indistinguishable from the pair of simple Pauli errors that constitute its detection events. In the example illustrated in Fig. 8, a Z error on the qubit marked with a circle between rounds ${r=3n}$ and ${r=3n+1}$ causes the same detection events as a combination of X and Y errors on that qubit occurring during the same time interval (or their indistinguishable counterparts). Let us label this qubit $q_{1}$ and consider a logical operator, $L$, with $q_{1}$ on its path. Between XX and YY measurements, $L$’s expression is either type one (X Paulis on blue checks) or type two (Y Paulis on green checks). The Z error affects $L$ only if its expression features X or Y Paulis on $q_{1}$. In such a case, one of the X or Y errors on $q_{1}$ anticommutes with $L$, while the other one commutes with it. As as result, a compound Pauli error and its indistinguishable simple Pauli error pairs impact logical operators in the same way, and the indistinguishability again does not pose an obstruction to error correction.
@@
The indistinguishability analysis also extends to measurement errors. A measurement error is indistinguishable from a combination of three simple Pauli errors. These equivalent Pauli errors for an erroneous measurement of a green check are shown in Fig. 9(b). The three Pauli errors consist of: first, a Z between ${r=3n-1}$ and ${r=3n}$; second, an X between ${r=3n}$ and ${r=3n+1}$; and third, a Y between ${r=3n+1}$ and ${r=3n+2}$. All these errors impact the physical qubit marked with a circle in Fig. 9(b). Equivalent Pauli errors for measurement errors of central blue and red checks can be obtained by $32\pi/$ 3 or $34\pi/$ 3 rotations in the hyperbolic plane, followed by one- or two-round shifts in time.
@@
A logical operator, $L$, with the faulty green edge on its closed path will incorporate the erroneous outcome of its measurement in the update protocol. This will lead to an incorrect sign in its expression. If $L$’s expression immediately before the measurement error at round $r=3n$ is type one (Z Paulis on green checks), its subsequent two expressions will be type two (Y Paulis on green checks) and, then, type one (Y Paulis on red checks). Among the three equivalent Pauli errors, the first and third one commute with $L$, while the second one anticommutes with it. Hence, their combined effect would be a sign flip in $L$’s expression, exactly as in the case of the measurement error.
@@
Analogously, if $L$’s initial expression immediately before the faulty measurement’s round is type two (X Paulis on red checks), its subsequent two expressions will be X Paulis on blue checks and Z Paulis on blue checks. Depending on weather $L$’s path traverses the red or blue check connected to the marked qubit, either the first or third Pauli error among the three equivalent errors will anticommute with $L$, while the other two will commute with it. Therefore, the combined effect of these equivalent Pauli errors is again the same as for the measurement error. Thus, the indistinguishability of measurement errors and combinations of simple Pauli errors does not pose an obstruction to error correction.
@@ Appendix B Threshold Plots with Various Duration of the Applied Errors
In Fig. 10 we present logical error rates of the hyperbolic Floquet codes under the entangling measurement error model for various duration of the applied errors. In all four panels the duration in which the errors are applied is proportional to the distance of the code so that the observed phase transition is a property of the three-dimensional space-time lattice of syndromes. To be more precise, this duration equals $hd$ rounds of check measurements with $h=1$-$4$ for panels (a)-(d). As seen here, the approximate value of the threshold error rate is around $0.25\,\%$ for all four panels and does not depend on the value of $h$. All plots in the main text are produced with $h=3$.
@@
Appendix C Honeycomb Floquet Codes from Symmetric Graphs

### v4

@@ Appendix A Details of Error Detection and Correction
In a measurement error, the outcome of a check’s measurement is faulty. This faulty outcome is then utilized in inferring two syndromes, causing them to exhibit detection events. An example is illustrated in Fig. 9(a). In this example, the highlighted red check’s outcome at round r=3n{r=3n} is erroneous, while all other measurements are accurate. This faulty red check leads to the green plaquette syndrome at round r=3n{r=3n} and the blue plaquette syndrome at round r=3n+1{r=3n+1} exhibiting detection events. Given the assumption that all other measurements are accurate, the next measurement of the same red check at round r=3n+3{r=3n+3} relates to the accurate value of the plaquette syndromes. Thus, the same green and blue syndromes generate additional detection events three rounds later. A measurement error, therefore, is linked to four detection events, forming a hyper-edge in the syndrome lattice, depicted with a dashed parallelogram in Fig. 9(a).
@@
Figure 9: Detection events of a measurement error. (a) The faulty measurement outcome of the highlighted red check at round r=3n{r=3n} is used to infer an incorrect value for the adjacent green and blue plaquette syndromes at rounds r=3n{r=3n} and r=3n+1{r=3n+1}. Later correct measurement of the same check leads to correct inference of these syndromes three rounds later. In both cases, the inferred syndrome differs from the most recent inference and gives rise to a detection event. Hence, a measurement error has four detection events forming a hyper-edge in the syndrome lattice, shown with a dashed parallelogram. (b) A combination of three simple Pauli errors on one of the qubits of the faulty check, denoted with a circle, gives rise to the same detection events. Three simple Pauli errors in this case are: a Z error between rounds r=3n1{r=3n-1} and r=3n{r=3n}, an X error between rounds r=3n{r=3n} and r=3n+1{r=3n+1}, and a Y error between rounds r=3n+1{r=3n+1} and r=3n+2{r=3n+2}. These equivalent Pauli errors have the same effects on the logical operators as the measurement error.
@@
Consider the pair of detection events linked to a simple Pauli error. There are four different Pauli errors that could trigger these two detection events and, thus, are indistinguishable from one another. Without any loss of generality, assume the simple Pauli error is an X error. The two plaquette syndromes with detection events correspond to adjacent green and blue octagons [see Fig. 8(b)]. These two octagons identify the location of the physical qubit affected by the X error to their shared red edge, without distinguishing which of the two ends of that edge it is. We label these two qubits as $q_{1}$ and $q_{2}$. Moreover, the X error could have occurred either between the most recent XX and YY measurements or between the most recent ZZ and XX measurements. These possibilities for error’s support qubit, $q_{1}$ or $q_{2}$, and its timing, between ZZ and XX measurements or between XX and YY measurements, give rise to the four indistinguishable errors with identical detection events.
@@
Now consider a logical operator, $L$, that contains one of these two qubits ($q_{1}$, without loss of generality) in its closed path. Between the most recent ZZ and XX measurements, $L$’s expression can either be type one (Z Paulis on red checks) or type two (X Paulis on blue checks). In each case, between the most recent XX and YY measurements, $L$’s expression will be type two (Y Paulis on red checks) or type one (X Paulis on green checks), respectively. The X error on $q_{1}$ affects $L$ only if $L$’s expression features Y or Z Paulis on $q_{1}$ at the time of the error. This holds true if $L$ is a type one and then a type two logical operator, in addition to having the red edge of $q_{1}$’s vertex along its closed path. In this case, $L$’s expression during the two possible time intervals includes Y or Z Paulis on both $q_{1}$ and $q_{2}$. Therefore, all four indistinguishable simple Pauli errors impact logical operators in the same way, and their indistinguishability does not pose an obstruction in correcting their effects on logical operators.
@@
A compound Pauli error is indistinguishable from the pair of simple Pauli errors that constitute its detection events. In the example illustrated in Fig. 8, a Z error on the qubit marked with a circle between rounds r=3n{r=3n} and r=3n+1{r=3n+1} causes the same detection events as a combination of X and Y errors on that qubit occurring during the same time interval (or their indistinguishable counterparts). Let us label this qubit $q_{1}$ and consider a logical operator, $L$, with $q_{1}$ on its path. Between XX and YY measurements, $L$’s expression is either type one (X Paulis on green checks) or type two (Y Paulis on red checks). The Z error affects $L$ only if its expression features X or Y Paulis on $q_{1}$. In such a case, one of the X or Y errors on $q_{1}$ anticommutes with $L$, while the other one commutes with it. As as result, a compound Pauli error and its indistinguishable simple Pauli error pairs impact logical operators in the same way, and the indistinguishability again does not pose an obstruction to error correction.
@@
The indistinguishability analysis also extends to measurement errors. A measurement error is indistinguishable from a combination of three simple Pauli errors. These equivalent Pauli errors for an erroneous measurement of a red check are shown in Fig. 9(b). The three Pauli errors consist of: first, a Z between r=3n1{r=3n-1} and r=3n{r=3n}; second, an X between r=3n{r=3n} and r=3n+1{r=3n+1}; and third, a Y between r=3n+1{r=3n+1} and r=3n+2{r=3n+2}. All these errors impact the physical qubit marked with a circle in Fig. 9(b). Equivalent Pauli errors for measurement errors of central green and blue checks can be obtained by $2\pi/3$ or $4\pi/3$ rotations in the hyperbolic plane, followed by one- or two-round shifts in time.
@@
A logical operator, $L$, with the faulty red edge on its closed path will incorporate the erroneous outcome of its measurement in the update protocol. This will lead to an incorrect sign in its expression. If $L$’s expression immediately before the measurement error at round $r=3n$ is type one (Z Paulis on red checks), its subsequent two expressions will be type two (Y Paulis on red checks) and, then, type one (Y Paulis on blue checks). Among the three equivalent Pauli errors, the first and third one commute with $L$, while the second one anticommutes with it. Hence, their combined effect would be a sign flip in $L$’s expression, exactly as in the case of the measurement error.
@@
Analogously, if $L$’s initial expression immediately before the faulty measurement’s round is type two (X Paulis on blue checks), its subsequent two expressions will be X Paulis on green checks and Z Paulis on green checks. Depending on whether $L$’s path traverses the blue or green check connected to the marked qubit, either the first or third Pauli error among the three equivalent errors will anticommute with $L$, while the other two will commute with it. Therefore, the combined effect of these equivalent Pauli errors is again the same as for the measurement error. Thus, the indistinguishability of measurement errors and combinations of simple Pauli errors does not pose an obstruction to error correction.
@@ Appendix B Threshold Plots with Various Duration of the Applied Errors
In Fig. 10 we present logical error rates of the hyperbolic Floquet codes under the entangling measurement error model for various duration of the applied errors. In all four panels the duration in which the errors are applied is proportional to the distance of the code so that the observed phase transition is a property of the three-dimensional space-time lattice of syndromes. To be more precise, this duration equals $hd$ rounds of check measurements with h=1h=1-44 for panels (a)-(d). As seen here, the approximate value of the threshold error rate is around $0.25\,\%$ for all four panels and does not depend on the value of $h$. All plots in the main text are produced with $h=3$.
@@
Appendix C Empirical Distances of Hyperbolic Floquet Codes
In this appendix, we investigate the performance of hyperbolic codes in suppressing logical errors. It is expected that at low physical error rates, $p$, the logical error rate of a code, $p_{L}$, is expected to decay exponentially with the code distance, dd[44]:
$p_{L}=A(d)p^{-\alpha d},$ (5)
where $A(d)$ is a coefficient dependent on the code distance and $\alpha$ is a constant.
The linear behavior of logical error rates in log-log plots under both error models (Figs. 11 and 12) provides numerical evidence for Eq. 5. In these figures, we also show linear fits for each line below the corresponding approximate error rate threshold. The extracted slopes from these fits should be proportional to what we call the empirical distances of the codes. Similar to the threshold rate, we expect empirical distances to depend on the code, the error model, and the decoder.
Figure 11: Any logical operator error rates for the family of hyperbolic Floquet codes under the phenomenological error model (dotted lines) [same as Fig. 4(a)] and linear fits to the values below the approximate error rate threshold of $p_{0}^{\rm all}\simeq 8\times 10^{-4}$ per round (solid lines).
Figure 12: Any logical operator error rates for the family of hyperbolic Floquet codes under the entangling measurements error model (dotted lines) [same as Fig. 5(a)] and linear fits to the values below the approximate error rate threshold of $p_{0}^{\rm all}\simeq 1.5\times 10^{-3}$ per round (solid lines).
Figure 13 shows the obtained empirical distances of hyperbolic codes under the phenomenological (panel a) and entangling measurements (panel b) error models. These empirical distances are assumed to be proportional to the slope of the fitted lines in Figs. 11 and 12. The proportionality coefficient is chosen such that, in each of the panels in Fig. 13, the square sum of the length of the dotted lines is minimized. These dotted lines connect each code’s data point to the solid line with a slope equal to 1. The closer a data point is to this line, the better the theoretical code distance predicts that code’s logical performance.
Whenever a code’s data-point falls to the left of the unit-slope solid line, the code is performing better in preserving logical qubits than predicted by its theoretical code distance. We anecdotally observe that while H16 and H2160 outperform their theoretical distances under both error models, and H144 and H400 underperform under both error models, the behavior of H64 changes with the error model. Specifically, H64 outperforms its theoretical code distance of $d=4$ under the phenomenological error model, with an empirical code distance closer to that of H144 ($d=6$), but underperforms under the entangling measurements error model, behaving similarly to H16 ($d=2$).
Figure 13: Empirical distances of hyperbolic codes under (a) phenomenological, and (b) entangling measurements error models. Theoretical code distances are defined as the lowest minimum weights of logical operators of a code (see Sec. 2). The empirical distance is proportional to the slope of a line fitted to the logical error rates of the codes below the approximate threshold value (see Figs. 11 and 12). If a code’s data point is to the left (right) of the diagonal line with a slope equal to 1, then it is outperforming (underperforming) compared to the average of the code family. For example, H64 outperforms its theoretical code distance of $d=4$ under the phenomenological error model but underperforms under the entangling measurements error model.
Appendix D Honeycomb Floquet Codes from Symmetric Graphs